# CallNavi: A Study and Challenge on Function Calling Routing and Invocation in Large Language Models

### v1

@@ Appendix A Meta Data
Appendix B Prompts
B.1 Model prediction
Prompt: Give the API list with descriptions below, then give the question in a chatbot, please give me the correct API that should be called. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Question} =======Question end======= Given the user question, and the APIs, classify and give a correct API name to call. The answer should be formatted in only one line, and only API names in brackets look like "[’getCustomerDetails’, ’depositFunds’]" or "[’getCustomerDetails’]". NO explanation and NO parameters in the answer!"
Prompt: Give the API list with descriptions below, then give the question in a chatbot, please give me the correct API that should be called. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Questions} =======Question end======= Given the user question, and the APIs, classify and give a correct API name and parameters to call. The answer should be formatted including API names and parameters in JSON style, which looks like: {’API’: [’getCustomerDetails’, ’depositFunds’], ’parameters’:[{"parameter1ForCall1": "***" },{"parameter1ForCall2": "***", "parameter2ForCall2": "***"}]} If we cannot get some parameter information from the question, set these parameters to "\$\$\$". NO explanation/notes in the answer!
The first prompt is designed for API calling list generating only in the experiment ’API Calling Exact Match’ from Table 6.1. The second prompt is related to generating full JSON output like ground truth in Section A.
B.2 LLM-as-judge
We use the following prompt to generate GPT Score of evaluation in Section 5.2
Prompt: You will be given a predicted API calling JSON and ground truth couple. Your task is to provide a ’total rating’ scoring on how similar the predicted API calling JSON is to ground truth JSON. Give your answer as a float on a scale of 0 to 1 with 3 decimals, where 0 means that the predicted JSON is not related to ground truth, and 1.0 means that the answer is completely the same as ground truth. You should consider if the API calling list is the same or not, and whether all parameters match or not. Only give a score in [[]], looks like [[1.00]]: ======Begin predict JSON====== {predict JSON} ======End of predict JSON====== =====Begin of ground truth JSON===== {ground truth JSON} =====End of ground truth JSON===== Response:
B.3 Backward Thinking
Prompt: Give the API list of a chatbot system with descriptions below, then give the question in chatbot, please first give me the one API that can answer the final task in the question. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Questions} =======Question end======= Given the user question, and the APIs, classify and give a correct last one task API name and parameters to call. The answer should be formatted into the following JSON format:{format JSON} We will solve the question in the next steps. NO explanation in the answer! ”’
After initial the session, give the following prompt to get a full API calling JSON.
Prompt: Then please based on the answer before, try to find APIs to fill in the missing information needed to complete the final API call list from the previous answer. Please also check if the API list has solved all sub-questions in the question. If not, please add other APIs and parameters related to the question in the correct order. The answer should be formatted in the JSON format used in the previous talk. Check if we can get all input parameters from the questions or not, if we can’t please don’t put <!stop> after the API list. If you think all the information is completed to solve the question, please write a <!stop> after the full API list. NO explanation in the answer! Only give API calling JSON!
Appendix C Case Study
C.1 Hallucination
"API": ["getLostBaggageReport", "updateBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}, {}]
However, the ground truth was:
"API": ["getBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}]
In this case, the model hallucinated two API calls, "getLostBaggageReport" and "updateBaggageStatus", which were not part of the provided API list. This hallucination led the model to predict incorrect API calls, deviating from the expected function "getBaggageStatus". Although the model correctly captured the parameter baggageId: "BAG123", it introduced an unnecessary second parameter block as an empty dictionary, further reducing the accuracy of the output.
C.2 JSON Generation
"API": ["getCustomerCreditCards"],
"parameters": [{"customerID": "123456"}]
(Assuming that ATM cards are considered credit cards for this specific API)
The ground truth, however, was:
"API": ["getATMCardList"],
"parameters": [{"accountID": "123456"}]
In this case, the model incorrectly generated an API call for "getCustomerCreditCards" instead of the correct API "getATMCardList". Additionally, the model included an unwanted note—"(Assuming that ATM cards are considered credit cards for this specific API)"—which made the output non-compliant with JSON formatting, as this additional text was outside the structure of the JSON object.
Appendix D Stability Research
@@
The stability score is calculated in Section 5.3. To give a clear example of calculating the stability of the model’s outputs, we analyze the frequency distribution of the results obtained from 5 times runs. Let’s review the variable settings:
@@
  • Tie Situations ($F_{1}=F_{2}$): When the maximum frequency is equal to the second maximum frequency, it indicates a tie for the most frequent output. In such cases, the stability score is set to 0 to reflect the lack of a clear majority or consensus.
  • No Tie Situations ($F_{1}>F_{2}$): The numerator $F_{1}-F_{2}$ measures the dominance of the most frequent output over the second most frequent one. The denominator $N-F_{2}$ normalizes this difference relative to the total number of outputs excluding those of the second most frequent output. The resulting score ranges between 0 and 1, where higher values indicate greater stability.
@@
    • Results: All outputs are different (e.g., $[^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime},^{\prime}D^{\prime},^{\prime}E^{\prime}]$).
@@
      Reflects complete instability due to lack of any consensus.
@@
  • Stability Score of 1: Perfect stability; all outputs are identical.
  • Stability Score of 0: No stability; either all outputs are unique or there’s a tie for the most frequent output.
@@
Usage:
@@
Table 7: Stability Test Results
@@
Appendix E Example Analysis of Hard and Medium Questions
This section provides an analysis of examples from the CallNavi dataset, focusing on the complexities of medium and hard questions in the bank scenario. The examples illustrate how the tasks test the capabilities of large language models (LLMs) in handling sequential API calls and parameter dependencies.
E.1 Hard Question Example
Question: "I want to travel to Japan, how many Japanese Yen I can spend with my credit card? My customer ID is 123155."
This question involves reasoning across multiple steps and APIs:
  1. 1.
    The model must first retrieve the customer’s credit cards using the getCustomerCreditCards API.
  2. 2.
    From the retrieved cards, the model must fetch details about a specific credit card using the getCreditCardDetails API.
  3. 3.
    Finally, the model must compute the currency exchange rate using the getCurrencyExchangeRates API to determine the spending amount in Japanese Yen.
This task tests the model’s ability to maintain dependencies between APIs, as the output of one API (e.g., credit card number) serves as an input for subsequent calls.
Listing 1: Hard Question Example from the Bank Scenario.
"id": "ban081",
"question": [
"role": "user",
"content": "I want to travel to Japan, how many Japanese Yen I can spend with my credit card? my customer ID is 123155."
"ground_truth": {
"API": ["getCustomerCreditCards", "getCreditCardDetails", "getCurrencyExchangeRates"],
"parameters": [
{ "customerID": "123155" },
{ "creditCardNumber": "\$\$\$" },
{ "currencyPair": "\$\$\$" }
"difficulty": "hard"
E.2 Medium Question Example
Question: "Retrieve details of wire transfer WT987654 and cancel it."
This question involves two sequential API calls:
  1. 1.
    The model must retrieve the details of the wire transfer using the getWireTransferDetails API.
  2. 2.
    The same transfer ID must then be used to cancel the transaction with the cancelWireTransfer API.
This task evaluates the model’s ability to handle basic sequential tasks with consistent parameter reuse across multiple APIs.
Listing 2: Medium Question Example from the Bank Scenario.
"id": "ban069",
"question": [
"role": "user",
"content": "Retrieve details of wire transfer WT987654 and cancel it."
"ground_truth": {
"API": ["getWireTransferDetails", "cancelWireTransfer"],
"parameters": [
{"transferID": "WT987654"},
{"transferID": "WT987654"}
"difficulty": "medium"
E.3 Analysis of Model Challenges
  • Hard Question: The main challenge lies in maintaining dependencies across APIs and correctly passing intermediate results (e.g., credit card number and currency pair).
  • Medium Question: While simpler than the hard question, this task requires consistency in parameter reuse (e.g., the same transferID must be used across both API calls).
These examples illustrate the varying levels of complexity in the CallNavi dataset, testing the models’ abilities to reason through multi-step tasks and maintain logical consistency in API usage.
E.4 Impact of JSON and YAML on Model Performance
@@
  • YAML to YAML: Both input and output data are formatted in YAML.
  • JSON to JSON: Both input and output data are formatted in JSON.
@@
The results of these experiments are shown in Table 8. Models were evaluated on multiple metrics:
  • Syntax Accuracy: Measures the correctness of the generated syntax.
  • Structure Accuracy: Evaluates whether the hierarchical structure of the output matches the ground truth.
  • Task Performance: Evaluates model performance across easy, medium, and hard tasks.
  • Overall Accuracy: Aggregates performance across all task levels.
  • GPT Score: A qualitative metric derived from GPT-based evaluation.
Table 8: Compare JSON or YAML as input/output performance differences.
@@
Model Syntax Acc Structure Acc easy medium hard overall gpt
@@
4. JSON to YAML is Challenging. The JSON to YAML configuration showed decreased performance compared to JSON to JSON, particularly in Syntax Accuracy (e.g., 0.598 for LLAMA 3.1). This indicates that YAML’s complexity as an output format negatively affects model performance, even when the input format is simpler.
Key Observations
  • JSON is more robust for both input and output, offering higher overall performance across tasks.
  • YAML introduces significant challenges in both generation and interpretation, particularly in harder tasks.
  • Mixed configurations, particularly YAML to JSON, mitigate some YAML-related difficulties, highlighting JSON’s simplicity as an output format.
These results underscore the importance of selecting appropriate data formats in real-world applications, as structured formats like YAML can introduce unexpected complexities that affect performance. This study also highlights the potential for optimizing LLMs to better handle YAML and similarly complex formats.
Appendix F Error Analysis
To better understand the limitations of current models in API function-calling tasks, we performed an error analysis on the predictions from our benchmark. This section highlights three typical types of errors observed in the generated outputs: hallucination, unexpected formatting, and logical inconsistencies.
F.1 Hallucination
Hallucination occurs when the model generates API names that do not exist in the provided API candidate pool. This issue often arises in scenarios where the model misinterprets the task requirements or over-generalizes from similar examples. For instance:
  • Example: The model generates an API call to getExchangeRateDetails, whereas the correct API name is getCurrencyExchangeRates.
  • Impact: Such errors result in invalid function calls and demonstrate the model’s difficulty in grounding its outputs to the provided API list.
F.2 Unexpected Formatting
Unexpected formatting errors include outputs containing invalid elements, such as triple backticks (“‘‘‘‘“‘), redundant notes, or commentary text. These errors are common when models generate outputs designed to mimic human explanations but fail to adhere strictly to the expected JSON or YAML format. Examples include:
  • Example: The model outputs:
    ‘‘‘json
    "API": "getCustomerDetails",
    "parameters": { "customerID": "12345"
  • Impact: These formatting errors make the outputs syntactically incorrect and require post-processing to remove extraneous elements.
F.3 Logical Errors in Hard Questions
Logical errors are particularly prevalent in hard questions, where the task involves multiple dependent API calls or complex reasoning. These errors include incorrect sequencing of API calls, failure to propagate parameters correctly, or omitting necessary steps. For example:
  • Impact: Logical errors highlight the limitations of current models in handling multi-step reasoning and task dependencies.
F.4 Summary of Error Types
The observed errors reveal challenges in three key areas:
  • Grounding to provided resources (hallucination).
  • Adherence to strict output formats (unexpected formatting).
  • Complex reasoning and multi-step tasks (logical errors).
Addressing these issues will require improving model training data, refining evaluation metrics, and incorporating techniques like grounding and reasoning-specific optimization.

### v2

@@ 4. CallNavi Dataset
Table 2. CallNavi dataset domains, questions and difficulties statistics table.
{tblr}
cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=3c, cell17 = r=2, hline1,3-14 = -, hline2 = 4-6, column7 = r Domain & API
Functions Questions Difficulty Max Input
Tokens
Easy Medium Hard
Banking 91 115 70 28 17 6517
Shopping 81 65 41 17 7 5195
Logistics 46 65 40 17 8 3434
Aviation 48 80 44 24 12 3461
Healthcare 20 47 31 10 6 1788
Public Services 82 85 50 27 8 6249
Human Resources 20 35 21 13 1 1863
Hotel Industry 49 65 40 19 6 3811
Insurance 42 60 40 11 9 3452
Telecommunications 100 112 79 22 11 6374
Overall 579 729 456 188 85 6517
4.1. Dataset
 [^1] https://github.com/Etamin/CallNavi
The CallNavi dataset evaluates LLMs’ task routing and API calling capabilities across multiple domains. As shown in Table 2, it contains 729 questions of varying difficulty and API interaction complexity, along with 579 distinct API functions. Questions are categorized into easy, medium, and hard levels:
  • Easy(456 questions): Require a single API call to fulfill task.
  • Medium(188 questions): Involve multiple APIs within the same question, with all parameters provided in the context.
5. Metrics and Evaluation
5.1. API Parameters AST Match
Our study utilizes Abstract Syntax Tree (AST) evaluation to assess models’ ability to generate accurate JSON outputs for API calls. The format of the output JSON and parameters follows a structure similar to the BFCL dataset in Section 4  (Yan et al., 2024). We parse the generated JSON string into an object and compare each component with the ground truth, such as the API list and parameters.
  • AST Exact Match: Whether the entire parsed object, including its structure and content, is identical to the ground truth.
Figure 2. Example of Evaluation Pipeline
This multistep evaluation ensures a thorough assessment of the accuracy of API function calls and the structural integrity of the parameters, allowing for a granular analysis of the performance.
Figure 3. Pipeline of AST Match Score
5.2. LLM-as-a-Judge Evaluation
We also use GPT-4o language models to evaluate whether the generated JSON outputs correspond accurately to the ground truth (Zheng et al., 2023). This approach aims to observe if LLMs can perform such evaluation tasks with high precision. Using an LLM for this purpose, we assess its ability to compare and validate structured data, thereby determining its effectiveness in automating the evaluation process.
5.3. Stability Score
  • $N$: Total number of outputs (samples).
The stability score is calculated as:
Stability $\text{Stability Score}=\frac{F_{1}-F_{2}}{N-F_{2}}$
To ensure reliable comparisons and reduce errors, we preprocess the outputs by removing unnecessary spaces, newlines, and formatting inconsistencies, converting the text to lowercase, and stripping extraneous characters that could cause mismatches.
While ‘freq@topk‘ is often used to evaluate the performance of model stability, it does not capture the stability in LLM output. For example, if a model produces the sequence ”AABBC” across multiple runs, ‘freq@topk‘ might assign a high score of 0.4 because the most frequent token (”A”) appears 40% of the time. However, this sequence is unstable as no single output consistently dominates. In contrast, our stability score focuses on the dominance of the most frequent output, offering a better measure of a model’s reliability in structured tasks.
@@ 5.3. Stability Score
To give a clear example of calculating the stability of the model’s outputs, we analyze the frequency distribution of the results obtained from 5 times runs. Let’s review the variable settings:
@@
  • Tie Situations ($F_{1}=F_{2}$): When the maximum frequency is equal to the second maximum frequency, it indicates a tie for the most frequent output. The stability score is set to 0 to reflect neither majority nor consensus in such cases.
  • No Tie Situations ($F_{1}>F_{2}$): The numerator $F_{1}-F_{2}$ measures the dominance of the most frequent output over the second most frequent one. The denominator $N-F_{2}$ normalizes this difference relative to the total number of outputs excluding those of the second most frequent output. The resulting score ranges from 0 to 1; higher values indicate greater stability.
@@
    • Results: All unique outputs (e.g., $[^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime},^{\prime}D^{\prime},^{\prime}E^{\prime}]$).
@@
      Complete instability due to a lack of consensus.
@@
  • Stability Score of 1: Perfect stability; all outputs are same.
  • Stability Score of 0: No stability; either all outputs are unique, or there’s a tie for the most frequent output.
@@
To further evaluate the stability of the model’s outputs, we calculate the average Levenshtein distance between the first answer and each subsequent output(Zhang et al., 2017). We normalize Levenshtein distance using the following formula.
6. Experiments and Benchmark
6.1. Models
To evaluate the benchmark, we selected models based on their performance, architecture, and relevance to function-calling tasks. The selection criteria focused on general-purpose and fine-tuned models optimized for function calling or JSON generation, ensuring a well-rounded comparison between zero-shot and fine-tuned capabilities. Models like BART and traditional retrieval-based approaches were excluded as they lack the ability to select APIs from extensive lists, which is critical for the complexity of this task.
Table 3 organizes the selected models into four groups: commercial models, medium-large models (10B + parameters), small models (5B-10B parameters) and light models (with parameters below 5B).
Model Name Origin Size Context Limits
GPT-4o OpenAI (Achiam et al., 2023) N/A 128K
GPT-4o-mini OpenAI (Achiam et al., 2023) N/A 128K
Gemini 1.5 Flash Google (Reid et al., 2024) N/A 1M
LLaMA 3.1 Meta AI (Meta-AI, 2024a) 70B 128K
Command-R Command AI (Cohere, [n. d.]) 35B 128K
Gemma2 Google (Team et al., 2024) 27B 8K
Mistral-Small Mistral AI (Ai, 2024a) 22B 128K
Phi3 Microsoft (Abdin et al., 2024) 14B 128K
Mistral-Nemo Mistral AI (Ai, 2024b) 12B 1M
Gemma2 Google (Team et al., 2024) 9B 8K
LLaMA 3.1 Meta AI (Meta-AI, 2024a) 8B 128K
xLAM Salesforce (Zhang et al., 2024) 7B 4K
DeepSeek R1 DeepSeek-AI (Guo et al., 2025) 7B 128K
NemoTron-Mini NVIDIA (Paul, 2024) 4B 4K
Phi3.5 Microsoft (Trufinescu, 2024) 3B 128K
LLaMA 3.2 Meta AI (Meta-AI, 2024b) 3B 128K
NexusRaven Nexusflow.ai (Nexusflow.ai, 2023) 13B 16K
Gorilla Berkeley (Patil et al., 2023) 7B 4K
Table 3. Comparison of General Purpose LLMs.
We chose 2 fine-tuned Function Calling models for testing, which have top performance on the BFCL leaderboard: NexusRaven and Gorilla OpenFunctions v2. Then we found that some models cannot input long lists, e.g. Firefunction v2 (Garbacki and Chen, 2024).
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4, cell17 = r=2, cell18 = r=2, cell19 = c=5, cell114 = r=2, cell31 = r=3, cell61 = r=6, cell121 = r=4, cell161 = r=3, cell191 = r=2, cell211 = r=3, hline1,24 = -0.08em, hline2 = 3-6,9-13, hline3 = -, Category & Models API Calling Routing
Exact Match Syntax
Validity Structural
Accuracy API Calling with Parameters
AST Match Overall
GPT Score
Easy Medium Hard All Easy Medium Hard All Avg. Macro Avg.
Commercial
Large
General
Middle
DeepSeek R1 7B 0.250 0.271 0.082 0.235 0.902 0.117 0.129 0.042 0.047 0.097 0.073 0.289
Light
Phi3.5 3B 0.723 0.340 0.188 0.562 0.004 0.0 0.0 0.0 0.0 0.0 0.0 0.002
Table 4. Benchmark Results Table. Macro Avg. means the arithmetic mean of 3 difficulties.
6.2. Environment
6.3. Pipeline
Our evaluation pipeline begins with the creation of prompts based on two templates. The first template focuses on retrieving the API calling list, which corresponds to the ”API” list in the ground truth. This prompt instructs the model to identify which API calls should be used and in what order. The second template is designed to generate the full API calling JSON, including the parameters for each call.
6.4. Benchmarks Results
The results presented in Table 6.1 highlight the performance of various models in different aspects of API function calling, including API calling routing accuracy, syntax validity, structural accuracy, and API parameter match through AST evaluation. OpenAI’s models, GPT4o and GPT4o mini, consistently outperform the others, particularly in syntax validity (0.993 and 0.994, respectively) and overall GPT score (0.913 and 0.908). Both models also demonstrate strong structural accuracy and API parameter AST match, especially in easier tasks. Gemini 1.5 Flash follows these metrics closely.
Our analysis demonstrates that the Pearson correlation between the ”GPT Score” and the ”All Avg.” column in the ”Parameter AST Match” section is 0.934, with a p-value of 4.40e-08. This indicates a very strong positive correlation, suggesting that higher GPT scores are closely associated with better average AST match performance. The results in our table are closely aligned with those of the Berkeley Function Calling Leaderboard [^2] https://gorilla.cs.berkeley.edu/leaderboard.html, which assesses LLMs’ performance in API or function-calling tasks. In both evaluations, models like OpenAI’s GPT4o stand out for their high syntax validity and overall accuracy, as reflected in our table where GPT4o scores above 0.9 in both categories. This matches the leaderboard’s top models, which also excel in AST evaluations and execution accuracy. In contrast, fine-tuned models such as FireFunction V2 in our results show weaker performance in API calling accuracy and AST matching, a trend similarly observed with fine-tuned models like Gorilla OpenFunctions on the Berkeley leaderboard, particularly in more complex API scenarios. Both evaluations emphasize the challenges faced by fine-tuned models in handling complex function scenarios or multi-step API calls, highlighting the need for improvement in these areas.
6.4.1. Stability Test
In our stability experiments, we ran 5 times referring to the previous study (Tian et al., 2023), and the stability results are in Table 5. By these metrics results, we obtain a numerical comparison that reflects the stability differences between outputs.
@@ 6.4.1. Stability Test
As mentioned in Section 5.3, a higher Election Stability Score indicates greater absolute consistency in the model’s outputs across multiple runs. A high Levenshtein Stability Score means similar between the same input in text generation output. Commercial models also perform better in Table 5 below.
Table 5. Stability Test Results
@@
DeepSeek R1 7B 0.058 0.501
7. Zero Shot Improvement
7.1. Calling + Parameters 2 Steps Generation
Figure 4. 2-Steps Generation Pipeline
To address these challenges, we propose combining the strengths of general-purpose LLMs and fine-tuned models shown in Figure4. Specifically, a general LLM selects the relevant APIs based on the input prompt, leveraging its superior understanding in identifying appropriate API calls. These selected APIs are then provided to a fine-tuned/LAM model, which focuses on generating the correct API calls along with the necessary parameters. This sequential process allows the general LLM to efficiently handle API selection, while the fine-tuned model concentrates on accurately producing API calls and parameters within a more manageable context.
As demonstrated in Table 6, this combined approach with GPT-4o routing significantly improves performance. Separating the tasks of API selection and parameter generation enhances the models’ ability to handle complex API calling tasks more effectively.
Table 6. 2 Steps Generation results for LLMs
Models easy medium hard overall
Gorilla v27B 0.682 0.005 0.000 0.427
xLAM-fc-7B 0.714 0.462 0.188 0.588
Gemma2 0.723 0.457 0.164 0.589
llama3.1 0.714 0.457 0.164 0.584
mistral-small 0.728 0.436 0.294 0.602
mistral-nemo 0.712 0.308 0.141 0.541
phi3:14b 0.019 0.005 0.011 0.015
command-r 0.633 0.547 0.223 0.563
llama3.2 0.462 0.297 0.082 0.375
nemotron-mini 0.208 0.01 0.000 0.133
LLM w/ itself as as router Gemma2:27b 0.598 0.59 0.341 0.566
Mistral-small 0.684 0.382 0.235 0.554
Command-r 0.621 0.505 0.247 0.547
7.2. Backward Inference Thinking
To optimize the API selection and calling process, we implement a Backward Thinking approach, inspired by CauseJudger (He and Lu, 2024) and Reverse Chain (Zhang et al., 2023), as illustrated in Figure 5. This approach enables the model to construct a sequence of API calls more systematically by working backwards from the final goal rather than following a purely forward selection strategy.
The process follows these steps:
  1. (1)
    Identifying the Final API Call: The model first determines the ultimate API needed to answer the user’s query. This API must provide the final required information or action.
  2. (2)
    Checking Parameter Completeness: The model verifies whether all required parameters for the final API are available. If any essential information is missing, the model does not proceed with execution but instead considers the necessary steps to obtain the missing data.
  3. (3)
    Determining Supporting API Calls: If missing parameters are identified, the model searches for additional APIs that can retrieve the necessary data. These supporting API calls are planned in reverse order, ensuring that the final API call has all the required inputs.
  4. (4)
    Iterative Refinement: This process continues iteratively. Each newly identified API is analyzed for its own dependencies, ensuring that all required information is recursively retrieved before execution.
By breaking the task into smaller, dependency-aware steps, this method allows the model to effectively plan and execute multi-step API calls, improving accuracy in complex scenarios. As shown in Table 7, this approach yields a 30% improvement in hard-level API calling tasks. The backward inference mechanism significantly enhances the model’s ability to handle intricate, real-world API calling scenarios, reducing failure cases caused by missing or misordered API dependencies.
Table 7. Backward Thinking performance in High difficulty calling in GPT-4o and GPT-4o-mini.
Model API Calling Routing API Calling with Parameters
Original
Backward
Thinking
Original
Backward
Thinking
GPT4o 0.611 0.894 0.388 0.729
GPT4o mini 0.564 0.847 0.364 0.482
Figure 5. Backward Thinking Pipeline
8. Case Study
8.1. Insufficient Context Limit
We observe that models like xLAM and NemoTron-Mini, which have a 4K context limit, struggle with longer API calls in CallNavi, where some inputs exceed 6K tokens. This limitation leads to truncated inputs, causing incorrect API selection and missing parameters in multi-call sequences. While models with higher context limits generally perform better, we also find that context length alone does not guarantee success—models must still effectively manage dependencies and navigate complex API workflows. These findings highlight the need for both expanded context windows and improved structured reasoning in function-calling tasks.
8.2. Hallucination
"API": ["getLostBaggageReport", "updateBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}, {}]
However, the ground truth was:
"API": ["getBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}]
In this case, the model hallucinated two API calls, ”getLostBaggageReport” and ”updateBaggageStatus”, which were not part of the provided API list. This hallucination led the model to predict incorrect API calls, deviating from the expected function ”getBaggageStatus”. Although the model correctly captured the parameter baggageId: ”BAG123”, it introduced an unnecessary second parameter block as an empty dictionary, further reducing the accuracy of the output.
8.3. JSON Generation
"API": ["getCustomerCreditCards"],
"parameters": [{"customerID": "123456"}]
#(Assuming that ATM cards are considered credit cards for this specific API)
The ground truth, however, was:
"API": ["getATMCardList"],
"parameters": [{"accountID": "123456"}]
In this case, the model incorrectly generated an API call for ”getCustomerCreditCards” instead of the correct API ”getATMCardList”. Additionally, the model included an unwanted note— ”(Assuming that ATM cards are considered credit cards for this specific API)”—which made the output non-compliant with JSON formatting, as this additional text was outside the structure of the JSON object.
8.4. Logical Errors in Hard Questions
Logical errors are particularly prevalent in hard questions, where the task involves multiple dependent API calls or complex reasoning. These errors include incorrect sequencing of API calls, failure to propagate parameters correctly, or omitting necessary steps. e.g.:
  • Impact: Logical errors highlight the limitations of current models in handling multistep tasks’ dependency reasoning.
8.5. Impact of JSON and YAML on Model Performance
@@
  • YAML to YAML: Both input and output are YAML.
  • JSON to JSON: Both input and output are JSON.
@@
The results of these experiments are shown in Table 8.
Table 8. Compare JSON or YAML as input/output performance differences.
@@
Model Syntax Acc Structure Acc Easy Medium Hard Overall GPT-Score
@@
4. JSON to YAML is Challenging. The JSON to YAML configuration showed decreased performance compared to JSON to JSON, particularly in Syntax Accuracy (e.g., 0.598 for LLAMA 3.1). This indicates that YAML’s complexity as an output format negatively affects model performance.
9. Discussion and Conclusion
9.1. Discussion
Our dataset introduces significant challenges, particularly in medium and hard questions, where models must select APIs from a large pool and generate parameters in multi-step and nested contexts. This complexity highlights the limitations of fine-tuned models trained on smaller API sets and underscores the need for more diverse and robust training paradigms.
We observe distinct model behaviours in API routing and parameter JSON generation. GPT-4o excels in both tasks, while models like LLaMA 3.1 and Gemma2 perform well in API routing but struggle with parameter generation, making them suitable for routing-centric applications. In contrast, smaller models (<10B parameters) exhibit instability, often producing inconsistent or incomplete outputs, limiting their effectiveness in complex, multi-step scenarios.
Long-context processing remains a significant bottleneck. Although models with larger context windows better handle structured inputs, they still struggle with simultaneous logical inference and structured JSON generation. Our findings suggest that merely increasing context size does not fully resolve multi-step reasoning challenges, emphasizing the need for improved architectures and reasoning strategies such as Chain of Thought.
Despite advancements, no model, including GPT-4o, fully solves intricate API calling tasks, reinforcing the need for further research in LLM-driven function calling.
9.2. Conclusion
This work introduces CallNavi, a benchmark evaluating API function calling in LLMs across 500 APIs and 700 questions. We assess general-purpose and fine-tuned models, revealing key limitations in API selection, parameter generation, and multi-step reasoning.
To improve function calling accuracy, we propose 2-steps generation and backward inference, enhancing structured API selection. While larger models like GPT-4o perform well, they still struggle with long-context input processing, particularly in tasks requiring both logical inference and structured JSON generation. Models with <6K token limits often truncate inputs, leading to incomplete API calls and degraded performance.
Our findings contribute to the broader field of software engineering evaluation and assessment, particularly in automated API function that calls for AI-based software design, stability evaluation and structured reasoning. Future work should focus on improving LLM robustness in real-world deployments, integrating retrieval-augmented techniques, and expanding function-calling benchmarks to incorporate real-time constraints such as error handling, authentication, and API versioning.
Threats To Validity
Internal Validity
One key limitation is context length constraints, where models like xLAM and NemoTron(4K) struggle with inputs exceeding 6K tokens in CallNavi, leading to truncation and incomplete API calls. While models with longer context windows perform better, our results suggest that context size alone is insufficient without strong reasoning and structured generation capabilities.
The complexity and variability of CallNavi, particularly in multi-step and nested API tasks, pose additional challenges. Fine-tuned models, often trained on smaller API sets, may struggle to generalize. Additionally, LLM-as-a-judge introduces potential subjectivity in evaluation. Our optimization strategies like 2-steps generation and backward inference—improve multi-step API selection, but their effectiveness may vary across different architectures.
External Validity
While CallNavi spans 500+ APIs and 700+ questions across 10 domains, it does not cover all real-world constraints, such as authentication, error handling, or evolving API versions. Future LLMS with longer context, hybrid architectures, etc., may demonstrate different performance trends.
Additionally, real-world API integration involves challenges beyond our benchmark, such as network failures, rate limits, and dynamic tool adaptation. While our evaluation covers syntax validity, AST match, and stability, future extensions should explore live production-level testing to better assess real-world deployment challenges.
Acknowledgment
We thank our collaborator BGL BNP Paribas for their support. The FNR funded this research under grants NCER22/IS/16570468/NCERFT and BRIDGES2021/IS/16229163/LuxemBERT.